OpenStudy (anonymous):

what is an equation of the line in slope intercept form. m=3and the y intercept is (0,4) a. y=4x+3 b. y=-3x+4 c. y=4x-3 d. y=3x+4

OpenStudy (jdoe0001):

use the point-slope intercept form \(\large \begin{array}{llll} &&slope = m=3\\ x_1&&y_1\\ (0&,&4)\\ \hline\\ &&&y-y_1=m(x-x_1) \end{array}\) and solve for "y"

OpenStudy (anonymous):

@wolf1728

OpenStudy (wolf1728):

I guess you want my help. I can figure this one out.

OpenStudy (anonymous):

yes please

OpenStudy (anonymous):

it should be d correct

OpenStudy (wolf1728):

We start with solving for 'b' b = y - mx substituting the values of slope and the point (0, 4) b = 4 -3•0 b = 4 Solving the complete equation y = mx + b y = 3x + 4